OpenStudy (anonymous):

what is 1! + 2! + 3! + ... + n!

myininaya (myininaya):

in sigma notation? it is \[\sum_{i=1}^{n}i!\]

OpenStudy (anonymous):

this means you add the factorial up to the last value that u r given

OpenStudy (anonymous):

yeah thats understood. but is there a general formula to get through this?

OpenStudy (amistre64):

there is if you write one up for it

OpenStudy (anonymous):

can some one please help me with this? i'm desperate :(

OpenStudy (amistre64):

an = (an-1) + n! an-1 = (an-2) + (n-1)! an = (an-2) +(n-1)! + n! = (an-3) +(n-2)! +(n-1)! + n! = (an-r) +(n-(r-1))! +(n-(r-2))!+...+(n-2)! + (n-1)! +n! n-r = 1 when r = (n-1) an = (an-(n-1)) +(n-((n-1)-1))! +(n-((n-1)-2))!+... ...+(n-2)! + (n-1)! +n! = (a1) +0! +3!+... +(n-2)! + (n-1)! +n! i wonder where if any a missed it

OpenStudy (amistre64):

1 (1+ 2(1 +3(1 +4(1+5+... a1 = 1 = 1 a2 = 1 + 2 = 3 a3 = 1 +2(1+3) = 9 a4 = 1 +2(1+3(1+4))) = 33 a5 = 1 +2(1+3(1+4(1+5)))) = 153 seems to work if it can be generalized

OpenStudy (amistre64):

an = 1 +2(1+3(1+4(...1+(n-1)(1+n)))))

OpenStudy (amistre64):

an = 1 +2(1+3(1+4(...1+(n-1)(1+n))))) an = 1 +2(1+3(1+4(n^2) ) ) ) )

OpenStudy (anonymous):

Um, there isn't really any formula in the usual closed form sense for the sum of first n factorials.

OpenStudy (anonymous):

amistre64 you approach kindles new thoughts but the final anwser you gave isn't general formula. :/

OpenStudy (anonymous):

As i said, there isn't one.

OpenStudy (anonymous):

il y mal.